Question 3 • AZ-Common Core Algebra II A
There are 10 integers in the 20s out of 50, so each number has probability 10/50 = 0.2 of being in the 20s. The probability that none of the first 4 numbers is in the 20s is 0.8⁴ = 0.4096. Therefore, the probability that at least one is in the 20s is 1 − 0.4096 = 0.5904, choice D.
AI-written and checked against the verified answer.